OpenStudy (tester97):

What is the 10th term of the sequence 625, 125, 25, ...?

OpenStudy (tester97):

@*Insert Name Here* halp pls?

OpenStudy (*insert name here*):

Do you know what kind of sequence this is?

OpenStudy (tester97):

geometric? arithmetic?

OpenStudy (tester97):

idk?

OpenStudy (*insert name here*):

Pretty sure it might be geometric.. so we need to find the ratio now~

OpenStudy (tester97):

1/5 is ratio

OpenStudy (*insert name here*):

Now just use the formula then :3

OpenStudy (tester97):

i did :/ but i didnt get the right answer.... \[a_{n}=a_{1} r^{n-1}\]then i substituted in the numbers \[a _{10}=625*(\frac{ 1 }{ 5 } )^9\] then i plugged it into the calculator and got some weird number :\

OpenStudy (*insert name here*):

Judging by the ratio and such you probably should xD Not all math is nice and rainbows

OpenStudy (tester97):

These are my answer choices\[\frac{ 1 }{ 3,125 }~\frac{1}{625}~\frac{1}{1,250}~\frac{1}{15,625}\] and and i dont know what to choose D:

OpenStudy (*insert name here*):

What answer did you get?

OpenStudy (tester97):

\[3.2_{E}-4\]

OpenStudy (*insert name here*):

What calculator you using?

OpenStudy (tester97):

TI-84 silver edition

OpenStudy (*insert name here*):

Awesome, alright. So once you get the precise answer from that equation. Go to "MATH" and then go to and press "Frac." Press enter again once it appears on the screen. c:

OpenStudy (tester97):

OMG I LOVE YOU FOR TELLING ME THAT ^_^

OpenStudy (*insert name here*):

*Comes in handy

OpenStudy (*insert name here*):

:)

OpenStudy (tester97):

Fail xD

OpenStudy (*insert name here*):

Oh shush >.>